You have $16000 you want to invest for the next 32 years. You are offered an investment plan that will pay you 8 percent per year for the next 16 years and 12 percent per year for the last 16 years. How much will you have at the end of the 32 years?
Answer #1

This question requires application of basic time value of money function, according to which

FV = PV * (1 + r)n

Now, for this question,

For first 16 years, r = 8%, n = 16, PV = $16,000

FV = 16000 * 3.4259

FV = $54,815.08

for this question,

For next 16 years, r = 12%, n = 16, PV = $54,815.08

FV = 54,815.08 * 6.1304

FV = $336,038.03 ---> Answer